OpenStudy (anonymous):

I need help solving x^2-1/2-35/16=0

OpenStudy (anonymous):

i guess you forgot a x in -1/2... right??

OpenStudy (anonymous):

yes I did

OpenStudy (anonymous):

multiply thru by 16 16x^2 - 8x - 35 = 0 (4x + 5 )(4x- 7 ) = 0 x = -5/4 or 7/4

OpenStudy (anonymous):

thank you!!! so much,

OpenStudy (anonymous):

use the following formula: \[(-b \pm \sqrt{b ^{2}-4ac} )/(2ac)\] a=1, b=-1/2, c= -35/16 this method solves EVERY function of this type.. ;-)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I understand Im suppose to use that fomaula but how to get rid of the fraction? simon

OpenStudy (anonymous):

don't.. just insert the fraction in the formula.. you'll be able to simplify it later..

OpenStudy (anonymous):

does it need further explanation?

OpenStudy (anonymous):

yes cause its two ways you can do it but using quadric formula or completing the square

OpenStudy (anonymous):

you were right Guru I have the answer thanks!